OpenStudy (anonymous):

which of the following is equivalent to x^5y^2/ xy^2 when x does not equal 0 and y does not equal 0? a. x^6y^5 b.x^5y c.x^4y d.x^4

hartnn (hartnn):

x^5/x = x^4 y^2/y^2 =1 so whats x^4*1 = ?

OpenStudy (anonymous):

x^4

OpenStudy (anonymous):

thank you

hartnn (hartnn):

yup, thats the answer, u got what i did ?

OpenStudy (anonymous):

yes, I see , thank you for your help :)